# Non-Asymptotic Performance Analysis of DOA Estimation Based on Real-Valued Root-MUSIC

### v1

@@ Abstract
This paper presents a systematic theoretical performance analysis of the Real-Valued root-MUSIC (RV-root-MUSIC) algorithm under non-asymptotic conditions. However, RV-root-MUSIC suffers from the problem of estimation ambiguity for the mirror roots, therefore the conventional beamforming (CBF) technique is typically employed to filter out the mirror roots. Through the equivalent subspace based on the conjugate extension method and the equivalence of perturbations for both true roots and mirror roots , this paper provides a comprehensive investigation of three critical aspects: noise subspace perturbation, true root perturbation, and mirror root perturbation characteristics in the RV-root-MUSIC algorithm. The statistical model is established and the generalized expression of perturbation is further developed. The simulation results show the correctness and validity of the derived statistical characteristics. The results provide a solid theoretical foundation for optimizing the parameter selection of DOA estimation in practical applications, particularly in radar systems, communication networks, and intelligent sensing technologies.
Index Terms—  RV-root-MUSIC, Direction-of-Arrival (DOA) estimation, non-asymptotic performance analysis, perturbation-theoretic expressions
@@ 1 Introduction
Direction of Arrival (DOA) estimation is one of nonlinear parameter estimation problems. Within this domain, the accurate acquisition of target parameters remains a central challenge in signal processing research. High-precision parameter estimation serves as a critical enabler for target perception across diverse applications, including radar detection, wireless communications, sonar localization, and medical imaging[1, 2, 3, 4, 5]. However, inherent noise contamination during signal transmission and reception introduces unavoidable errors. Consequently, estimation accuracy emerges as a key metric for evaluating algorithmic performance. To develop robust error quantification frameworks and systematic performance analysis methodologies is significant for DOA estimation [6].
The Multiple Signal Classification (MUSIC)[7] algorithm has revolutionized DOA estimation by achieving super-resolution performance. However, its high computational complexity due to spectral peak search limits its practical applications. To address this, the root-MUSIC[8] algorithm was developed, significantly reducing computational load. Despite these advancements, MUSIC and its variants operate in the complex domain, which still poses efficiency challenges[9]. Consequently, real-valued subspace-based algorithms, such as Real-Valued MUSIC (RV-MUSIC) [10] and Real-Valued root-MUSIC (RV-root-MUSIC) [11, 12], have been proposed to further enhance computational efficiency. While the conventional root-MUSIC exhibits satisfactory performance under high signal-to-noise ratio (SNR) and asymptotic conditions, its estimation accuracy degrades significantly in low-SNR scenarios or limited snapshots (non-asymptotic conditions). Performance analysis of conventional complex-domain DOA estimation algorithms has been conducted[13, 14], while the performance of real-valued algorithms remains to be analysed.
This paper presents a non-asymptotic performance analysis of RV-root-MUSIC. Theoretical error expressions are derived for both true roots and mirror roots, with their validity confirmed through Monte Carlo simulations demonstrating agreement between theory and simulation.
@@ 2.1 The model of the signal and arrays
where $d$ denotes the inter-element spacing, and $\lambda$ represents the signal wavelength. $\bm{S}(t)\in\mathbb{C}^{K\times M}$ is the source signal matrix.
@@ 2.2 Review about RV-root-MUSIC
Based on the defined noise subspace, the root-finding polynomial for DOA estimation algorithm can be constructed as:
@@
As demonstrated in [11], RV-root-MUSIC suffers from the mirror roots, where the relationship $\theta_{mirror}=-\theta_{true}$ is satisfied. Thus the conventional beamforming (CBF) technique is typically employed to filter out the mirror roots.
@@ 3 Non-asymptotic Performance Analysis of RV-root-MUSIC
This paper makes a statistical characteristic analysis of real roots and mirror roots about RV-root-MUSIC under non-progressive conditions and also verifies the estimation accuracy with respect to the number of array elements.
@@ 3.1 Deviation of the true DOAs
Given that $\bm{\mathbb{E}}^{T}_{n}\bm{p}(r_{k})=0$ and the relation of (17), (16) can be simplified and we can further obtain the perturbation expression of $\bm{\mathbb{E}}_{n}$:
@@ 3.2 Deviation of the mirror DOAs
$\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}^{*}}\doteq 2jAr_{k}\mathrm{Im}(\frac{\Delta r_{k}^{*}}{r_{k}^{*}})K_{t}(r_{k}^{*})G(r_{k}^{*})$ (21)
@@ 4 Statistical Properties of True and Mirror DOA Perturbations
Substituting (30) into (20) and (24), identical forms can be observed. To simplify performance analysis model, the general perturbation expression $\Delta_{k}$ can be expressed as:
@@ 5 Simulation Experiments and Results
condition 1: Varying the SNR from $0dB$ to $20dB$ in $2dB$ steps with $200$ snapshot, and the result is shown in Fig. 2.
@@
6 Summary
This paper formulated a framework of the non-asymptotic performance analysis of the RV-root-MUSIC algorithm, deeply analyzed the perturbation characteristics of the noise subspace, the true roots and the mirror roots, established a statistical model, and conducted simulation experiments. The final results show that the research in this paper can provide a tighter theoretical lower bound for the performance of the RV-root-MUSIC algorithm. In addition, the study on the parity of the array elements in this paper provides guidance for the performance improvement of the algorithm: that is, choosing an odd number of array elements to ensure the pairing phenomenon of the roots and thereby enhance the robustness of the algorithm.
7 Acknowledgment
The work of X. T. Meng is supported by the Shandong Provincial Natural Science Foundation (No.ZR2024QF068) and the Aeronautical Science Foundation of China (No.2023Z037077002). The work of F. Gini and M. Greco has been partially supported by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).

### v2

@@ Abstract
This paper presents a systematic theoretical performance analysis of the Real-Valued root-MUSIC (RV-root-MUSIC) algorithm under non-asymptotic conditions. A well-known limitation of RV-root-MUSIC is the estimation ambiguity caused by mirror roots, which are typically suppressed using conventional beamforming (CBF). By leveraging the equivalent subspace constructed through the conjugate extension method and exploiting the equivalence of perturbations for true and mirror roots, this work provides a comprehensive study of three key aspects: noise subspace perturbation, true-root perturbation, and mirror-root perturbation. A statistical model is established, and generalized perturbation expressions are derived. Monte Carlo simulations confirm the correctness and effectiveness of the theoretical results. The analysis provides a rigorous foundation for parameter optimization in Direction-of-Arrival (DOA) estimation, with applications in radar, wireless communications, and intelligent sensing.
Index Terms—  RV-root-MUSIC, DOA estimation, non-asymptotic performance analysis, perturbation-theoretic expressions
@@ 1 Introduction
Direction-of-Arrival (DOA) estimation is a fundamental nonlinear parameter estimation problem. Accurate parameter acquisition is essential for reliable target perception in diverse applications, including radar detection, wireless communications, sonar localization, and medical imaging [1, 2, 3, 4, 5, 6, 7]. However, unavoidable noise contamination during signal transmission and reception introduces estimation errors, making accuracy a key criterion for evaluating algorithmic performance. Developing robust error quantification frameworks and systematic performance analysis methodologies is therefore of significant importance [8].
The Multiple Signal Classification (MUSIC) algorithm [9] marked a breakthrough in DOA estimation by achieving super-resolution performance. However, its reliance on exhaustive spectral peak searches results in high computational complexity, which limits practical applicability. To address this issue, the root-MUSIC algorithm [10] was introduced, significantly reducing computational burden. Nonetheless, both MUSIC and root-MUSIC operate in the complex domain, which can still pose efficiency challenges [11]. To further improve efficiency, real-valued subspace-based approaches such as Real-Valued MUSIC (RV-MUSIC) [12] and Real-Valued root-MUSIC (RV-root-MUSIC) [13, 14] have been proposed. While root-MUSIC performs well under high signal-to-noise ratio (SNR) and asymptotic conditions, its accuracy deteriorates considerably in low-SNR or limited-snapshot scenarios, i.e., non-asymptotic conditions. Although performance analyses for conventional complex-domain DOA estimation algorithms exist [15, 16], real-valued methods have not been thoroughly examined.
This paper addresses this gap by presenting a non-asymptotic performance analysis of RV-root-MUSIC. Theoretical error expressions are derived for both true and mirror roots, and their validity is verified through Monte Carlo simulations, which demonstrate close agreement between theoretical predictions and empirical results.
@@ 2.1 The model of the signal and arrays
where $d$ denotes the inter-element spacing, and $\lambda$ is the signal wavelength. $\bm{S}(t)\in\mathbb{C}^{K\times M}$ is the source signal matrix.
@@ 2.2 Review about RV-root-MUSIC
Based on the defined noise subspace, the root-finding polynomial for RV-root-MUSIC can be constructed as:
@@
As demonstrated in [13], RV-root-MUSIC suffers from the mirror roots, where the relationship $\theta_{mirror}=-\theta_{true}$ is satisfied. Thus the conventional beamforming (CBF) technique is typically employed to filter out the mirror roots as below:
$\left|\bm{a}^{H}(\theta)\bm{R}_{xx}\bm{a}(\theta)\right|^{2}$ (8)
@@ 3 Non-asymptotic Performance Analysis of RV-root-MUSIC
This paper makes a statistical characteristic analysis of real roots and mirror roots about RV-root-MUSIC under non-asymptotic conditions and also verifies the estimation accuracy with respect to the number of array elements.
@@ 3.1 Deviation of the true DOAs
Given that $\bm{\mathbb{E}}^{T}_{n}\bm{p}(r_{k})=0$ and the relation of (18), (17) can be simplified and we can further obtain the perturbation expression of $\bm{\mathbb{E}}_{n}$ as (19):
@@ 3.2 Deviation of the mirror DOAs
$\begin{split}\left.\frac{\partial f(z,\tilde{r})}{\partial z}\right|_{z=r_{k}^{*}}&=A[K_{m}(r_{k}^{*})^{{}^{\prime}}[K_{t}(r_{k}^{*})G(r_{k}^{*})]\\ &\quad\left.+K_{m}(r_{k}^{*})[K_{t}(r_{k}^{*})G(r_{k}^{*})]^{{}^{\prime}}]\right|_{\tilde{r}=r_{k}^{*}+\Delta(r_{k}^{*})}\\ &=\left.A[K_{m}(r_{k}^{*})^{{}^{\prime}}K_{t}(r_{k}^{*})G(r_{k}^{*})]\right|_{\tilde{r}=r_{k}^{*}+\Delta(r_{k}^{*})}\\ &\doteq 2jAr_{k}Im(\frac{\Delta r_{k}^{*}}{r_{k}^{*}})K_{t}(r_{k}^{*})G(r_{k}^{*})\end{split}$ (22)
@@
[…]
Given that $\bm{\mathbb{E}}^{T}_{n}\bm{p}(r_{k}^{*})=0$ and the relation of (24), (23) can be simplified and we can further obtain the perturbation expression of $\bm{\mathbb{E}}_{n}$ as (25):
@@ 4 Statistical Properties of True and Mirror DOA Perturbations
Substituting (33) into (21) and (27), identical forms can be observed. To simplify the performance analysis model, the general perturbation expression $\Delta_{k}$ can be expressed as:
@@ 5.1 Experimental Setup and Results
condition 1: Varying the SNR from $0dB$ to $20dB$ in $2dB$ steps with $200$ snapshots, and the result is shown in Fig. 2.
@@
From Fig. 2, the RMSE decreases as the SNR increases, and the simulation performance aligns with the theoretical performance more and more closely.
@@
From Fig. 3, the RMSE decreases with increasing snapshot number, though the rate of improvement diminishes progressively. Beyond a certain threshold, further increases in computational complexity and processing time yield marginal performance gains. Consequently, selecting an appropriate snapshot count that balances resolution accuracy and computational efficiency is critical.
@@
5.2 Experimental Conclusion
From the above experiments, it can be seen that the trends of the simulation performance and theoretical performance varying with the SNR and the number of snapshots are basically consistent. The theoretical performance provides a tighter theoretical lower bound for RV-root-MUSIC.
6 Conclusion
This paper presented a non-asymptotic performance analysis framework for the RV-root-MUSIC algorithm. Unlike prior works that mainly focus on asymptotic or complex-domain analyses, the proposed framework systematically characterizes the perturbation behavior of the noise subspace, true roots, and mirror roots in the real-valued setting. A statistical model was developed, and generalized perturbation expressions were derived, providing a unified formulation for both true and mirror DOA deviations. Monte Carlo simulations confirmed the theoretical predictions, demonstrating close alignment with the analytical results.
Beyond the theoretical contributions, this work offers several practical insights:
  • The analysis establishes a tighter theoretical performance bound for RV-root-MUSIC in finite-snapshot and low-SNR regimes, which are highly relevant to real-world radar, communications, and sensing systems.
  • The investigation of array element parity revealed that even-element ULAs suffer from root degeneracy on the real axis, leading to spurious DOAs at broadside directions. The proposed correction factor and the recommendation to employ odd-element arrays provide a design guideline for practitioners to enhance algorithm robustness.
  • The derived perturbation expressions enable parameter optimization (e.g., array size, number of snapshots, and SNR trade-offs), offering practical value for system design under constrained computational resources.
Overall, this paper advances the theoretical understanding of RV-root-MUSIC under non-asymptotic conditions and bridges the gap between abstract perturbation theory and practical DOA estimation design. Future work will extend the framework to correlated sources, arbitrary array geometries, and adaptive mirror-root suppression techniques, further enhancing its applicability in modern sensing and communication platforms.
7 Funding Acknowledgements
The work of X. T. Meng is supported by the Shandong Provincial Natural Science Foundation (No.ZR2024QF068). The work of F. Gini and M. Greco has been partially supported by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).